An investment offers a 12% total return over the coming year. Bill Morneau thinks the total real return on this investment will be only 7%.

What does Morneau believe the inflation rate will be over the next year? (Do not round intermediate calculations. Round the final answer to 2 decimal places.)
Answer #1

Solution :

As per the Fisher Equation formula

( 1 + Nominal Rate ) = ( 1 + Real Rate ) * ( 1 + Inflation Rate )

As per the information given in the question we have

Total Return = Nominal Rate = 12 % = 0.12

Real rate = 7 % = 0.07

Inflation rate = To find

Applying the above information in the fisher equation formula we have

( 1 + 0.12 ) = ( 1 + 0.07 ) * ( 1 + Inflation rate )

( 1.12 ) = ( 1.07 ) * ( 1 + Inflation rate )

1.12 / 1.07 = ( 1 + Inflation rate )

1.046729 = ( 1 + Inflation rate )

( 1 + Inflation rate ) = 1.046729

Inflation rate = 1.046729 – 1

Inflation rate = 0.046729

Inflation rate = 4.6729 %

Inflation rate = 4.67 % ( when rounded off to two decimal places )

Thus the Inflation rate over the next year = 4.67 %
Answer #2

Wesimann Co. issued 14-year bonds a year ago at a coupon rate of 7.4 percent. The bonds make semiannual payments and have a par value of $1,000. If the YTM on these bonds is 5.7 percent, what is the current bond price?
Answer #3

ANSWER :


Let I be inflation rate.

Real rate = (1 + Nominal rate) / ( 1 + I)  - 1

=> I = (1 + Nominal rate) / (1 + real rate) - 1

=> I = (1 + 0.12) / (1 + 0.07) - 1

=> I = 0.0467

=> I = inflation rate =  4.67% (ANSWER).

answered by: Tulsiram Garg